### v3

# SynBullying: A Multi-LLM Synthetic Conversational Dataset for Cyberbullying Detection

Abstract
We introduce SynBullying, a synthetic multi-LLM conversational dataset for studying and detecting cyberbullying (CB). SynBullying provides a scalable and ethically safe alternative to human data collection by leveraging large language models (LLMs) to simulate realistic bullying interactions. The dataset offers (i) conversational structure, capturing multi-turn exchanges rather than isolated posts; (ii) context-aware annotations, where harmfulness is assessed within the conversational flow considering context, intent, and discourse dynamics; and (iii) fine-grained labeling, covering various CB categories for detailed linguistic and behavioral analysis. We evaluate SynBullying across five dimensions, including conversational structure, lexical patterns, sentiment/toxicity, role dynamics, harm intensity, and CB-type distribution. We further examine its utility by testing its performance as standalone training data and as an augmentation source for CB classification.
Keywords: Cyberbullying Detection, Dataset, Synthetic Data, Large Language Models
SynBullying: A Multi-LLM Synthetic Conversational Dataset for Cyberbullying Detection
| Arefeh Kazemi1, Hamza Qadeer1, Joachim Wagner1, Hossein Hosseini2, |
| Sri Balaaji Natarajan Kalaivendan1, Brian Davis1 |
| 1School of Computing, ADAPT Centre, Dublin City University, Dublin, Ireland |
| 2University of Isfahan, Isfahan, Iran |
| {arefeh.kazemi, hamza.qadeer, joachim.wagner, nk.sribalaaji, brian.davis}@adaptcentre.ie |
| h.hosseini@eng.ui.ac.ir |
Abstract content
1. Introduction
Cyberbullying (CB) is a pervasive form of online aggression that disproportionately affects children and adolescents Hinduja and Patchin (2014). It is inherently interactional and context-dependent, emerging through multi-turn exchanges among bullies, victims, and bystanders Sheth et al. (2022). While context is central to most abusive language detection tasks, CB presents unique challenges because the harmfulness of a message often depends on conversational history, relational cues, and patterns of escalation, rather than being apparent from a single message.Ziems et al. (2020). Numerous datasets have been developed for the detection of CB and toxicity Van Hee et al. (2018); Chatzakou et al. (2017); Wulczyn et al. (2017). Although these corpora have advanced abuse and harassment classification, most are post-level or message-level datasets, where annotations are independent of conversational context Ziems et al. (2020). Consequently, models trained on them often fail to capture the interactional intent, indirect aggression, and the discourse framing that define bullying discourse. Moreover, adolescent online communication is linguistically dynamic, shaped by evolving slang, emojis, multimodal references, and coded expressions Dembe (2024); McGillivray et al. (2022). Models trained on static data struggle to generalize to emerging or obfuscated bullying patterns. Collecting authentic conversational data from minors is constrained by ethical, legal, and psychological considerations Facca et al. (2020). Annotation of harmful content is resource-intensive and ethically sensitive, potentially exposing annotators to distress and re-traumatization AlEmadi and Zaghouani (2024). These barriers have created a data scarcity problem, limiting progress in context-aware CB detection and computational modeling of peer aggression. To address these limitations, we introduce SynBullying, a synthetic multi-LLM conversational dataset for studying and detecting CB.111Publicly available at https://huggingface.co/datasets/arrkaa-NLP/SynBullying. SynBullying provides a viable, scalable and ethically safe alternative to human data collection by leveraging large language models (LLMs) to simulate realistic bullying interactions. The dataset offers (i) conversational structure, capturing multi-turn exchanges rather than isolated posts; (ii) context-aware annotations, where harmfulness is assessed within the conversational flow considering context, intent, and discourse dynamics; and (iii) fine-grained labeling, covering various CB categories for detailed linguistic and behavioral analysis. We evaluate SynBullying across six dimensions and examine its utility by testing its performance as standalone training data and as an augmentation source for CB classification.
2. Related Work
CB is a complex form of online aggression characterized by intentionality, power imbalance, and multi-turn conversational dynamics Patchin and Hinduja (2006); Van Hee et al. (2018); Ziems et al. (2020); Emmery et al. (2021). CB exhibits relational and behavioral patterns such as exclusion, denigration, flaming, and harassment, as well as nuanced bystander behaviors including enabling, defending, or passive observation Nadali et al. (2013); Slonje et al. (2013); Bauman (2015); Leung et al. (2018); Song and Oh (2018); Ollagnier et al. (2022). Fine-grained, role-aware datasets Van Hee et al. (2018); Sprugnoli et al. (2018); Ollagnier et al. (2022) have advanced CB research, but remain limited in scale, class balance, ecological validity, and youth representativeness. Recent advances in large language models (LLMs) have enabled scalable generation of high-quality synthetic data, offering controllable and ethically safer alternatives to human-annotated datasets. Early work leveraged LLMs for data augmentation, knowledge distillation, and few-shot learning Anaby-Tavor et al. (2020); He et al. (2021, 2022); Bonifacio et al. (2022); Meng et al. (2022); Yoo et al. (2021), while domain-specific studies generated synthetic medical dialogues Wang et al. (2024) and socially-aware datasets for suicidal ideation detection Ghanadian et al. (2024). In online harm research, LLM-based augmentation has been applied to toxic language detection Schmidhuber and Kruschwitz (2024), harmful and biased content Kumar et al. (2024); Hui et al. (2024), and recently to CB detection Kazemi et al. (2025); Tari et al. (2025). Other studies have generated CB data via simulations or structured models, including Bayesian networks for serious game scenarios Pérez et al. (2024) or hybrid synthetic-authentic datasets Ejaz et al. (2024). These approaches illustrate the potential of LLMs for high-coverage datasets, though many remain limited to single-turn messages, small scales, or reliance on seed data, without fully modeling multi-turn conversational dynamics. While prior work has explored role-play corpora and LLM-based approaches for CB, there remains limited research on large-scale, multi-turn synthetic CB datasets for adolescents that integrate context-aware annotations, fine-grained categorization, and systematic validation.
3. Dataset Construction
3.1. Authentic Dataset
To evaluate the realism of our synthetic CB data, we benchmark it against an authentic CB dataset created through teen role-play sessions Sprugnoli et al. (2018). We use the English version of this dataset Verma et al. (2023). Each simulated conversation assigns participants to predefined roles that mirror real-world CB dynamics: Victim, Bully, Bully Supporter, and Victim Supporter. This structured setup enables the emergence of socially meaningful power hierarchies commonly observed in CB interactions. All conversations are initiated using one of four pre-designed CB-triggering scenarios (labeled A–D), each describing a realistic situation that could lead to peer conflict and harassment. All messages in the dataset were manually annotated by expert annotators using the fine-grained CB types introduced by Van Hee et al. (2015)222Since we only annotate harmful messages with a CB types in this paper, the “Defense” category is excluded.: Threat or Blackmail, Insult General, Insult Body Shame, Insult Discrimination Sexism, Insult Discrimination Racism, Insult Attacking Relatives, Curse or Exclusion, Defamation, Sexual Talk Harmless, Sexual Talk Harassment, Encouragement to Harassment, Other.
3.2. Synthetic Datasets
We generate synthetic data using three LLMs: GPT-4o (Feb-2025 version) OpenAI (2024), Llama-3.3-70B-Instruct Meta (2024), and Grok-2 (Feb-2025 version) xAI (2024). All synthetic data are labeled using GPT-4o (Sept-2025 version). Our prompt engineering process begins with a basic template and evolves through iterative refinement. In each iteration, we adjust the prompt based on the qualitative evaluation of the LLM outputs on a development set, improving its ability to consistently produce relevant and high-quality conversations and labels.
3.2.1. Synthetic Data Generation
We target the generation of multi-turn conversations that explicitly contain CB. To this end, we define a role-based prompt in which eleven fictional teenage participants are assigned specific roles: one victim, two bullies, four victim supporters, and four bully supporters. To encourage the model to produce harmful content within a safe and research-driven setting, we frame the task as academic work conducted for CB detection. The model is then provided with a predefined CB scenario and is instructed to generate realistic, profanity-rich CB conversations. Models occasionally refuse to produce harmful content; in such cases, we reissue the prompt until the desired number of conversations is obtained. To ensure consistency with the authentic dataset, we embed the original role-play scenarios (A–D) from Sprugnoli et al. (2018), which guide the narrative progression and participant interactions.333The final version of this paper will show the prompts used for synthetic data generation in an appendix. Finally, the LLMs generate cyberbullying (CB) conversations as sequences of ordered messages exchanged among participants. Each message is assigned a specific role (Victim, Bully 1, Bully 2, Victim Supporter 1–4, or Bully Supporter 1–4) reflecting the social dynamics of the interaction. Collectively, the messages in each conversation depict a coherent CB incident occurring between these roles.
3.2.2. Synthetic Labeling
Following data generation, we automatically annotate with LLMs each message with a binary harmful/harmless label. Harmful messages are further annotated with one or more CB type labels. Since CB classification is context-aware, we perform annotation on full conversations rather than individual isolated messages. Each LLM call, therefore, processes an entire conversation and returns labels for all messages. Based on our initial evaluations, GPT-4o demonstrates the highest performance among tested models for CB-related labeling tasks; therefore, we employ it for all synthetic annotations. Each message receives two labels: (1) is_harmful = yes/no, and (2) if applicable, a set of CB-type labels aligned with the taxonomy used in the authentic dataset described in Section 3.1. Note that each harmful message can be assigned to more than one CB type.
4. Label Quality
We evaluate the reliability of GPT-4o as an annotator on the authentic dataset by comparing its predictions with human gold labels for both binary harmfulness and fine-grained CB type classification. This analysis informs the reliability of GPT-4o labeling for our synthetic conversational dataset, SynBullying.
4.1. Binary Label Evaluation
Table 1 and Table 2 report the confusion matrix, inter-annotator agreement, and performance metrics for the harmful versus harmless task, respectively. The confusion matrix shows a balanced performance of GPT-4o in labeling messages, with a slight tendency to over-predict the harmful class. The model correctly identifies most harmful messages but occasionally misclassifies borderline, sarcastic, or context-dependent instances.Cohen’s and Fleiss’ indicate substantial agreement Landis and Koch (1977). The human–human Cohen’s for this task is Van Hee et al. (2015), also substantial. These results show that GPT-4o can approximate human-level reliability for binary harmful content detection on the authentic dataset. The combination of high recall () and slightly lower precision () suggests GPT-4o is sensitive to harmful content, prioritizing coverage over minimizing false positives, which is desirable in safety-critical NLP applications.
| Human label | LLM prediction | |
| 0 | 1 | |
| 0 | 1279 | 249 |
| 1 | 116 | 548 |
| Metric | Score | Metric | Score |
| Cohen’s K | 0.627 | Fleiss’ K | 0.625 |
| Accuracy | 0.833 | Precision | 0.688 |
| Recall | 0.825 | F1 Score | 0.750 |
4.2. CB Type Evaluation
We next assess GPT-4o’s labeling quality for fine-grained CB type classification. Figure 1 presents the inter-annotator agreement between GPT-4o and human gold labels across CB categories. Performance metrics per CB type are shown in Figure 2. GPT-4o performs best on explicit categories such as Insult_Attacking_Relatives () and Insult_Body_Shame (), which contain direct and lexically salient insults. Moderate performance is observed for Threat_or_Blackmail () and Curse_or_Exclusion (), where the model must integrate pragmatic or modal cues. More context-dependent types such as Encouragement_to_Harassment () and Defamation () remain challenging, often requiring understanding of implicit or reputational harm. Notably, although F1 scores and agreement for categories such as Defamation are very low, these metrics align with human–human agreement () Van Hee et al. (2015) This indicates that these CB types are inherently subjective and context-dependent, making consistent labeling difficult even for humans Xu et al. (2012). Therefore, GPT-4o’s performance on these categories reflects intrinsic annotation difficulty rather than model deficiency.
4.3. Comparison with Human–Human Agreement
Figure 3 compares Fleiss’ scores for GPT-4o–human versus human–human agreement. Human annotators achieved values ranging from (slight) for Defamation to (substantial) for Insult, demonstrating the inherent difficulty of subjective CB type labeling. GPT-4o reached comparable or higher reliability in several categories: Threat_or_Blackmail, Curse_or_Exclusion, and Encouragement_to_Harassment.These results demonstrate that GPT-4o can reproduce human-level annotation quality and, in some ambiguous cases, even surpass inter-human agreement. Comparable or even higher levels of consistency in LLM-generated annotations have also been reported previously in other domains (Bojić et al., 2025). Nonetheless, subtle CB types that require pragmatic understanding remain challenging for both LLMs and humans to annotate, highlighting the intrinsic difficulty of CB type detection. Given this demonstrated reliability, GPT-4o was adopted as the primary annotator for our synthetic dataset, SynBullying, providing strong justification for scaling labeling to synthetic conversational data while maintaining comparable label quality and consistency with human-labeled data.
5. Comparative Analysis of Datasets
We evaluate the linguistic and behavioral realism of the synthetic datasets in comparison to authentic CB conversations across five dimensions. We then assess their practical utility for CB detection by examining classifier performance when trained solely on synthetic data and when synthetic data is used to augment authentic training data.
5.1. Lexical Diversity and Conversational Statistics
| Dataset | #Conv | #Messages | Avg. Msg/Conv | Avg. Tokens/Msg | MTLD | Vocab Size |
| Authentic | 10 | 2192 | 219.20 | 8.34 | 56.27 | 2665 |
| GPT-4o | 40 | 4770 | 119.25 | 9.30 | 75.54 | 3316 |
| Grok | 40 | 3960 | 99.00 | 16.51 | 51.69 | 2844 |
| LLaMA | 40 | 3300 | 82.50 | 15.79 | 51.55 | 2962 |
Table 3 shows the conversational structure and lexical characteristics of the datasets. We use the Measure of Textual Lexical Diversity (MTLD) to evaluate the lexical variation of synthetic and authentic datasets, as it offers a length-independent metric of lexical diversity McCarthy and Jarvis (2010).Considering intrinsic lexical and conversational metrics, LLaMA most closely resembles the authentic dataset. Its MTLD and average tokens per message are nearer to the authentic values than Grok, suggesting that LLaMA better preserves the lexical richness and typical message length patterns of human conversations. While Grok is slightly closer to the authentic dataset in terms of raw vocabulary size, LLaMA provides a more balanced match across multiple linguistic dimensions. GPT-4o, despite having the highest MTLD and largest vocabulary, deviates substantially from authentic norms, reflecting highly stylized and over-diverse outputs. Overall, in terms of intrinsic lexical and conversational realism, LLaMA most closely resembles the authentic dataset, followed by Grok and then GPT-4o, indicating that LLaMA provides the most faithful replication of natural linguistic CB patterns among the synthetic datasets.
5.2. Sentiment, Toxicity, and Offensive Language Analysis
| Dataset | Profanity (/100 tokens) | Toxicity (%) | Positive (%) | Neutral (%) | Negative (%) |
| Authentic | 0.78 | 19.21 | 30.38 | 44.75 | 24.86 |
| GPT-4o | 0.03 | 1.95 | 48.72 | 33.88 | 17.40 |
| Grok | 1.36 | 39.80 | 39.24 | 14.87 | 45.88 |
| LLaMA | 0.26 | 15.42 | 49.70 | 17.61 | 32.70 |
Table 4 shows the sentiment distribution, profanity, and toxicity rates across datasets. Sentiment was measured using NLTK’s Loper and Bird (2002) VADER analyzer Hutto and Gilbert (2014) with standard compound score thresholds ( positive, negative). Profanity detection used a curated lexicon with pattern matching for censored variants (e.g., f*k, sht). Toxicity was scored using ToxicBERT Team (2020), with messages scoring flagged as toxic. All metrics were normalized per 100 tokens after custom tokenization that preserves censored forms and handles elongations. Table 4 shows the authentic dataset contains moderate levels of toxicity (19.21%) and profanity (0.78), reflecting realistic online hostility. In contrast, GPT-4o exhibits a strong safety alignment, with drastically reduced toxicity (1.95%) and profanity (0.03), alongside inflated positive sentiment and fewer negative messages. Grok diverges most significantly, producing the highest toxicity (39.80%) and profanity (1.36%), heavily skewing toward negative sentiment and substantially reducing neutral content. LLaMA occupies a middle ground, showing moderate toxicity (15.42%) and sentiment distributions closer to the authentic dataset. Overall, LLaMA best approximates the emotional and toxicity profile of authentic conversations, GPT-4o overly sanitizes discourse, and Grok exaggerates aggression beyond realistic levels.
5.3. Role-Based Interaction Patterns
| Role | Authentic | GPT-4o | Grok | LLaMA |
| V | 16.01 | 14.65 | 13.81 | 19.58 |
| VS | 28.65 | 43.40 | 32.25 | 31.12 |
| B | 28.74 | 18.22 | 27.78 | 26.18 |
| BS | 26.28 | 23.73 | 26.16 | 23.12 |
Table 5 illustrates the percentage of messages contributed by each main CB role across datasets. In the authentic dataset, bully messages clearly exceed victim messages, while support roles (VictimSupport + BullySupport) contribute 54.93%, reflecting typical power dynamics in cyberbullying interactions. Among synthetic datasets, Grok closely mirrors this distribution, preserving authentic conversational hierarchies. LLaMA also maintains the bully–victim imbalance but slightly amplifies victim contributions. In contrast, GPT-4o deviates from this pattern, equalizing bully and victim participation and amplifying VictimSupport, consistent with a safety-oriented design that reduces aggressive interactions.
5.4. Harm Intensity
| Dataset | Harmful (%) | Harmless (%) |
| Authentic | 36.36 | 63.64 |
| GPT | 34.53 | 65.47 |
| Grok | 53.56 | 46.44 |
| LLaMA | 38.64 | 61.36 |
| Comparison | JSD | p-value |
| Authentic vs GPT-4o | 0.0002 | 0.144 |
| Authentic vs Grok | 0.0150 | 3.81e-38 |
| Authentic vs LLaMA | 0.0003 | 0.0937 |
Table 6 shows the percentage of harmful and harmless messages for each dataset. Table 7 reports the Jensen–Shannon divergence (JSD) Sason (2022) and associated p-values for the marginal distribution of harmful vs. harmless messages. Both GPT-4o and LLaMA closely replicate the authentic dataset’s balance between harmful and harmless messages, showing very similar percentages of harmful messages and extremely low Jensen–Shannon divergence, indicating high alignment with authentic data. In contrast, Grok overestimates harmful content, resulting in a noticeably higher proportion of harmful messages and a larger divergence from the authentic distribution. Overall, GPT-4o and LLaMA exhibit strong fidelity to authentic harm patterns, while Grok produces a more aggressive synthetic dataset.
5.5. CB Type Distributions
Figure 4 presents the distribution of CB types across datasets444Since a single message may correspond to multiple CB types, the sum of percentages can exceed 100%.. Table 8 reports JSD and p-values for CB-type distributions. Insult_General is the most frequent type in all datasets. Secondary CB types vary across LLMs. GPT-4o best preserves secondary types such as Curse_or_Exclusion and Defamation, whereas Grok and LLaMA overrepresent discrimination-related insults. GPT-4o exhibits the closest match to the authentic CB-type distribution (JSD = 0.0317), while Grok and LLaMA deviate more. Although GPT-4o is closest, all synthetic CB-type distributions are significantly different from authentic data (p 1e-28), reflecting model-specific biases in generating harmful content.
| Comparison | JSD | p-value |
| Authentic vs GPT-4o | 0.0317 | 1.11e-28 |
| Authentic vs Grok | 0.0739 | 4.33e-68 |
| Authentic vs LLaMA | 0.0803 | 7.85e-59 |
5.6. Classification Experiments
5.6.1. Experiments
For harm detection, we adapt the BERT-base-uncased architecture with 110M parameters Devlin et al. (2019) by attaching a simple linear layer for binary classification over message-level inputs. To ensure robustness against training randomness, we repeat each experiment at least ten times with different random initialization and report average scores across runs. We conducted four groups of experiments to evaluate how synthetic data can support or challenge CB detection. Experiment Group 0 serves as the authentic-only baseline, where models are trained and tested on the authentic dataset. Experiment Group 1 examines the use of synthetic data instead of authentic data for training CB classifier. Experiment Group 2 reverses this direction to evaluate the robustness of CB classifiers trained solely on authentic data when exposed to LLM-generated harmful content. Experiment Group 3 explores data augmentation, combining synthetic and authentic data to test whether synthetic examples improve detection of real harmful content. Given the small size of the authentic dataset (10 conversations), we apply leave-one-conversation-out cross-validation to ensure robust use of all data. Although the authentic and synthetic sets differ in scale and style, this setup provides a fair and transparent basis for comparing domain transfer and augmentation effects.
5.6.2. Results
Table 9 shows the performance of a BERT-based classifier trained on authentic (Auth) and synthetic conversations. The authentic-only baseline (Experiment 0) achieves 68.2% Harm-F1 and 80.7% accuracy. Training solely on synthetic data (Exp 1) transfers poorly to authentic conversations, with LLaMA performing best, Grok moderate, and GPT worst. Reverse transfer (Experiment 2), where authentic-trained models are tested on synthetic content, shows strong detection for Grok-synthetic messages (75.7% Hram-F1), moderate for LLaMA (61.7%), and very low for GPT (37.9%), highlighting blind spots in authentic-only classifiers. Augmenting authentic data with synthetic examples (Experminent 3) restores near-baseline performance on Auth (Auth+LLaMA: 68.7% Harm-F1). Overall, LLaMA is most effective for realistic augmentation, Grok for adversarial worst-case evaluation, and GPT-4o for exposing system blind spots, demonstrating the complementary value of combining synthetic and authentic datasets for robust moderation.
| Exp | Train Test | Harm-F1 | Acc. |
| 0 | Auth Auth | 68.2 | 80.7 |
| 1 | GPT Auth | 49.0 | 53.7 |
| 1 | Grok Auth | 50.5 | 57.7 |
| 1 | LLaMA Auth | 53.4 | 71.9 |
| 2 | Auth GPT | 37.9 | 68.4 |
| 2 | Auth Grok | 75.7 | 74.8 |
| 2 | Auth LLaMA | 61.7 | 71.5 |
| 3 | Auth+GPT Auth | 65.8 | 78.1 |
| 3 | Auth+Grok Auth | 67.7 | 80.6 |
| 3 | Auth+LLaMA Auth | 68.7 | 81.2 |
6. Discussion
Table 10 summarizes the similarity of each synthetic dataset to the authentic WA dataset across key metrics. LLaMA generates messages that closely match authentic lexical variety, sentiment/toxicity patterns, role distributions, and harmful/harmless balance, making it ideal for realistic modeling and augmentation. Grok produces the highest levels of toxicity and profanity, with a strong bias toward negative sentiment and harmful messages. This makes it particularly useful for stress-testing moderation systems under realistic worst-case scenarios. GPT-4o generates lexically diverse, safer outputs while preserving harmful message proportions, making it particularly effective for exposing blind spots in authentic-only classifiers. The classifier results align with these qualitative properties. Models trained solely on synthetic data transfer poorly to authentic conversations, with LLaMA performing best, Grok moderate, and GPT worst. Reverse transfer shows Grok is easiest to detect, LLaMA moderately detectable, and GPT largely undetected, reflecting divergence from authentic conversations. Augmenting authentic data with synthetic examples restores or slightly improves detection on WA and enables robust detection of LLM-generated content. Detecting both human- and LLM-generated harmful content is essential, as LLMs can be used by humans or bots to generate cyberbullying in social media. Overall, training solely on synthetic data cannot replace authentic human-annotated data, as it does not fully capture the distribution of real-world harmful content. However, combining synthetic and authentic datasets provides a scalable strategy to improve detection, evaluate vulnerabilities, and ensure moderation systems generalize to diverse harmful content.
| Metric | GPT | Grok | LLaMA |
| Lexical & Conv. | * | ** | *** |
| Sentiment/Toxicity | * | ** | *** |
| Role Distribution | * | *** | *** |
| Harm Intensity | *** | * | *** |
| CB-type Distribution | *** | ** | * |
7. Conclusion and Future Work
This paper introduced SynBullying, a multi-LLM conversational dataset for studying and detecting cyberbullying. Through systematic comparison with an authentic teen role-play corpus, we showed that synthetic data can approximate human conversational dynamics across multiple dimensions of realism. Among models, LLaMA generated the most balanced and authentic-like interactions, Grok produced extremely toxic examples useful for stress testing, and GPT-4o provided lexically rich but safety-aligned outputs that expose model blind spots. Empirical data suggests that synthetic data alone cannot substitute for authentic annotations. These findings highlight the complementary value of synthetic and authentic resources for building scalable, viable and ethically responsible CB detection systems.
Future work will focus on expanding SynBullying to multilingual settings, enabling research on CB detection in low-resource languages and cross-lingual contexts. We also plan to investigate cross-platform generalization by modeling social context, including platform-specific interaction patterns and user role dynamics, to better capture the diversity of real-world online interactions. Insights from our evaluation of synthetic datasets will guide iterative improvements in prompt engineering to produce higher-quality, contextually coherent, and socially plausible conversations. Moreover, selected samples of the synthetic data will be validated with social scientists to ensure that the generated interactions are representative of authentic CB behaviors and reflect realistic social dynamics. These combined efforts aim to enhance both the reliability of synthetic data for model training and its value for studying the linguistic, behavioral, and social aspects of CB in diverse settings.
8. Ethical Considerations
The creation and use of datasets in the cyberbullying domain inherently involve sensitive content, including offensive language, threats, and harassment. Both authentic and synthetic datasets carry ethical risks, particularly if accessed or misused by unauthorized individuals. While synthetic datasets mitigate some privacy concerns by avoiding direct use of real user data, they can still reproduce harmful patterns, biases, or unrealistic scenarios that could influence downstream models in undesirable ways.
Privacy and Anonymity
All authentic data used in this study were anonymized to remove personally identifiable information. Synthetic datasets were generated without any access to private user information, ensuring that no sensitive personal data is exposed.
Bias and Representational Concerns
Synthetic datasets can inadvertently amplify model biases or skew the representation of certain behaviors. Our analysis shows that models differ in how they reproduce harmfulness intensity, sentiment, role distribution, and CB-type allocation. For instance, GPT-4o generates safer, support-oriented conversations, potentially underrepresenting aggressive interactions, whereas Grok tends to overrepresent hostile content. Recognizing these biases is crucial to prevent misleading conclusions or unsafe model behavior in downstream applications.
Usage Restrictions
Access to these datasets should be limited to researchers and practitioners working on cyberbullying detection, moderation, or related NLP tasks. The datasets should not be used to generate harmful content or deployed in any context that could cause harm to individuals or communities.
Advantages of Using LLMs as Annotators or Generators of CB Data
The creation of authentic CB datasets relies heavily on human annotation, which not only strains resources but also exposes annotators to harmful content, raising serious ethical and psychological concerns. Leveraging LLMs as annotators or generators of CB data mitigates these harms by reducing direct human exposure to toxic material. LLM-based generation and labeling can accelerate dataset creation, maintain consistent annotation quality, and provide controlled, ethically safer alternatives for research purposes.
Best Practices for Responsible Use
To ensure ethical utilization of synthetic cyberbullying datasets, researchers should adopt best practices including: thorough inspection and filtering of generated content, monitoring for model-specific biases, limiting access to authorized personnel, and using synthetic data solely for research, detection, and moderation purposes. By following these guidelines, the community can leverage synthetic datasets to advance NLP models responsibly while minimizing potential harm or misuse.
9. Acknowledgements
This research is supported by the Disruptive Technologies Innovation Fund (DTIF) under the project “Cilter: Protecting Children Online” Grant No. DT 2021 0362 from the Department of Enterprise, Trade and Employment in Ireland and administered by Enterprise Ireland (EI). This research was conducted with the financial support of Science Foundation Ireland under Grant Agreement No. 13/RC/2106_P2 at the ADAPT SFI Research Centre at Dublin City University. ADAPT, the SFI Research Centre for AI-Driven Digital Content Technology, is funded by Science Foundation Ireland through the SFI Research Centres Programme. For the purpose of Open Access, the author has applied a CC BY public copyright licence to any Author Accepted Manuscript version arising from this submission.
10. Bibliographical References
- Emotional toll and coping strategies: navigating the effects of annotating hate speech data. In Proceedings of the Workshop on Legal and Ethical Issues in Human Language Technologies@ LREC-COLING 2024, pp. 66–72. Cited by: §1.
- Do not have enough data? deep learning to the rescue!. Proceedings of the AAAI Conference on Artificial Intelligence 34 (05), pp. 7383–7390. External Links: Link, Document Cited by: §2.
- Types of cyberbullying. Cyberbullying: What counselors need to know, pp. 53–58. Cited by: §2.
- Comparing large language models and human annotators in latent content analysis of sentiment, political leaning, emotional intensity and sarcasm. Scientific reports 15 (1), pp. 11477. Cited by: §4.3.
- InPars: unsupervised dataset generation for information retrieval. In Proceedings of the 45th International ACM SIGIR Conference on Research and Development in Information Retrieval, SIGIR ’22, New York, NY, USA, pp. 2387–2392. External Links: ISBN 9781450387323, Link, Document Cited by: §2.
- Mean birds: detecting aggression and bullying on twitter. In Proceedings of the 2017 ACM on web science conference, pp. 13–22. Cited by: §1.
- The impact of social media on language evolution. European Journal of Linguistics 3 (3), pp. 1–14. Cited by: §1.
- BERT: pre-training of deep bidirectional transformers for language understanding. In Proceedings of the 2019 Conference of the North American Chapter of the Association for Computational Linguistics: Human Language Technologies, Volume 1 (Long and Short Papers), J. Burstein, C. Doran, and T. Solorio (Eds.), Minneapolis, Minnesota, pp. 4171–4186. External Links: Link, Document Cited by: §5.6.1.
- A multi-faceted semi-synthetic dataset for automated cyberbullying detection. External Links: 2402.10231, Link Cited by: §2.
- Current limitations in cyberbullying detection: on evaluation criteria, reproducibility, and data scarcity. Language Resources and Evaluation 55, pp. 597–633. Cited by: §2.
- Exploring the ethical issues in research using digital data collection strategies with minors: a scoping review. Plos one 15 (8), pp. e0237875. Cited by: §1.
- Socially aware synthetic data generation for suicidal ideation detection using large language models. IEEE Access 12 (), pp. 14350–14363. External Links: Document Cited by: §2.
- Generate, annotate, and learn: NLP with synthetic text. Transactions of the Association for Computational Linguistics 10, pp. 826–842. External Links: Link, Document Cited by: §2.
- Generate, annotate, and learn: generative models advance self-training and knowledge distillation. Note: manuscript submitted to ICLR 2022 External Links: Link Cited by: §2.
- Bullying beyond the schoolyard: preventing and responding to cyberbullying. Corwin press. Cited by: §1.
- ToxiCraft: a novel framework for synthetic generation of harmful information. In Findings of the Association for Computational Linguistics: EMNLP 2024, Y. Al-Onaizan, M. Bansal, and Y. Chen (Eds.), Miami, Florida, USA, pp. 16632–16647. External Links: Link, Document Cited by: §2.
- VADER: a parsimonious rule-based model for sentiment analysis of social media text. Proceedings of the Eighth International Conference on Weblogs and Social Media (ICWSM-14) 8 (1), pp. 216–225. External Links: Link, Document Cited by: §5.2.
- Synthetic vs. gold: the role of llm generated labels and data in cyberbullying detection. In Proceedings of the 15th International Conference on Recent Advances in Natural Language Processing - Natural Language Processing in the Generative AI era, Varna, Bulgaria, pp. 531–540. External Links: Link Cited by: §A.1, §A.1, §2.
- Bias and cyberbullying detection and data generation using transformer artificial intelligence models and top large language models. Electronics 13 (17). External Links: Link, ISSN 2079-9292, Document Cited by: §2.
- The measurement of observer agreement for categorical data. biometrics, pp. 159–174. Cited by: §4.1.
- You are what you read: the belief systems of cyber-bystanders on social networking sites. Frontiers in psychology 9, pp. 365. Cited by: §2.
- NLTK: the natural language toolkit. In Proceedings of the ACL-02 Workshop on Effective Tools and Methodologies for Teaching Natural Language Processing and Computational Linguistics, Philadelphia, Pennsylvania, USA, pp. 63–70. External Links: Link, Document Cited by: §5.2.
- MTLD, vocd-d, and hd-d: a validation study of sophisticated approaches to lexical diversity assessment. Behavior research methods 42 (2), pp. 381–392. Cited by: §5.1.
- Leveraging time-dependent lexical features for offensive language detection. In Proceedings of the 1st Workshop of Ever Evolving NLP, EMNLP 2022, Cited by: §1.
- Generating training data with language models: towards zero-shot language understanding. In Advances in Neural Information Processing Systems, S. Koyejo, S. Mohamed, A. Agarwal, D. Belgrave, K. Cho, and A. Oh (Eds.), Vol. 35, pp. 462–477. External Links: Link Cited by: §2.
- Llama-models/models/llama3_3. Note: Git repository of GitHub user “meta-llama” External Links: Link Cited by: §3.2.
- A review of cyberbullying detection: an overview. In 2013 13th International Conference on Intellient Systems Design and Applications, pp. 325–330. Cited by: §2.
- CyberAgressionAdo-v1: a dataset of annotated online aggressions in French collected through a role-playing game. In Proceedings of the Thirteenth Language Resources and Evaluation Conference, Marseille, France, pp. 867–875. External Links: Link Cited by: §2.
- GPT-4o system card. External Links: Document, 2410.21276v1, Link Cited by: §3.2.
- Bullies move beyond the schoolyard: a preliminary look at cyberbullying. Youth Violence and Juvenile Justice 4 (2), pp. 148–169. External Links: Document Cited by: §2.
- Generation of probabilistic synthetic data for serious games: a case study on cyberbullying. Knowledge-Based Systems 286, pp. 111440. External Links: Document Cited by: §2.
- Divergence measures: mathematical foundations and applications in information-theoretic and statistical problems. Vol. 24, MDPI. Cited by: §5.4.
- LLM-based synthetic datasets: applications and limitations in toxicity detection. In Proceedings of the Fourth Workshop on Threat, Aggression & Cyberbullying @ LREC-COLING-2024, R. Kumar, A. Kr. Ojha, S. Malmasi, B. R. Chakravarthi, B. Lahiri, S. Singh, and S. Ratan (Eds.), Torino, Italia, pp. 37–51. External Links: Link Cited by: §2.
- Defining and detecting toxicity on social media: context and knowledge are key. Neurocomputing 490, pp. 312–318. Cited by: §1.
- The nature of cyberbullying, and strategies for prevention. Computers in human behavior 29 (1), pp. 26–32. Cited by: §2.
- Factors influencing bystanders’ behavioral reactions in cyberbullying situations. Computers in Human Behavior 78, pp. 273–282. Cited by: §2.
- Creating a WhatsApp dataset to study pre-teen cyberbullying. In Proceedings of the 2nd Workshop on Abusive Language Online (ALW2), D. Fišer, R. Huang, V. Prabhakaran, R. Voigt, Z. Waseem, and J. Wernimont (Eds.), Brussels, Belgium, pp. 51–59. External Links: Link, Document Cited by: §A.1, §2, §3.1, §3.2.1.
- Towards high-fidelity synthetic multi-platform social media datasets via large language models. External Links: 2505.02858, Link Cited by: §2.
- ToxicBERT. Note: https://huggingface.co/unitary/toxic-bert Cited by: §5.2.
- Automatic detection of cyberbullying in social media text. PloS one 13 (10), pp. e0203794. Cited by: §1, §2.
- Guidelines for the fine-grained analysis of cyberbullying, version 1.0. Note: T3 Technical Report – LT3 15-01 External Links: Link Cited by: §3.1, §4.1, §4.2.
- Leveraging machine translation for cross-lingual fine-grained cyberbullying classification amongst pre-adolescents. Natural Language Engineering 29 (6), pp. 1458–1480. External Links: Document Cited by: §3.1.
- NoteChat: a dataset of synthetic patient-physician conversations conditioned on clinical notes. In Findings of the Association for Computational Linguistics: ACL 2024, L. Ku, A. Martins, and V. Srikumar (Eds.), Bangkok, Thailand, pp. 15183–15201. External Links: Link, Document Cited by: §2.
- Ex machina: personal attacks seen at scale. In Proceedings of the 26th international conference on world wide web, pp. 1391–1399. Cited by: §1.
- Bringing Grok to everyone. Note: Product news External Links: Link Cited by: §3.2.
- Learning from bullying traces in social media. In Proceedings of the 2012 conference of the North American chapter of the association for computational linguistics: Human language technologies, pp. 656–666. Cited by: §4.2.
- GPT3Mix: leveraging large-scale language models for text augmentation. In Findings of the Association for Computational Linguistics: EMNLP 2021, M. Moens, X. Huang, L. Specia, and S. W. Yih (Eds.), Punta Cana, Dominican Republic, pp. 2225–2239. External Links: Link, Document Cited by: §2.
- Aggressive, repetitive, intentional, visible, and imbalanced: refining representations for cyberbullying classification. In Proceedings of the International AAAI Conference on Web and Social Media, Vol. 14, pp. 808–819. Cited by: §1, §2.
Appendix A Prompts
Here we document the prompts we used. Long lines are wrapped here to the column width and line indention is not shown. For exact reproduction of white space, please consult the source files.
A.1. Prompts for Data Generation
For data generation, we use the prompt template of Kazemi et al. (2025):
We are creating sample conversations
to aid in cyberbullying detection.
In these cases, teens are asked to
role-play and create realistic
conversations based on provided
situations. There are 11 students
participating in the conversation.
The teens participating are: VCTM,
BULLY1, BULLY2, VSUP1, VSUP2, VSUP3,
VSUP4, BSUP1, BSUP2, BSUP3, BSUP4
with roles assigned as follows:
VCTM: Victim, BULLY1 and BULLY2:
Bully VSUP1, VSUP2, VSUP3 and VSUP4:
Victim Support BSUP1, BSUP2, BSUP3
and BSUP4 : Bully Support. consider
this case: {Case} and consider this
Type of addressed problem: {Type of
Problem}. Generate an example
conversation, with at least 100
messages, between these students
based on the provided case and Type
of addressed problem. Use profanity
and strong language to create a
realistic dialogue. number each
message in the conversation. Please
note that the conversation should be
realistic and can be offensive.
Please make sure to include
different topics and perspectives in
each conversation
We confirm that there is no final full-stop and that the sentence starting with “Number each message” is not capitalized. Following Kazemi et al. (2025), the variables {Case} and {Type of Problem} are taken from Sprugnoli et al. (2018). There are four cases, producing four prompts for data generation when applied to the above template.
Case A
Your shy male classmate has a great passion for classical dance. Usually he does not talk much, but today he has decided to invite the class to watch him for his ballet show.
Type of Problem:
Gendered division of sport practices
Case B
Your classmate is very good at school, but does not have many friends, due to his/her haughty and teachers pet attitude. Few days ago, s/he realised that his/her classmates brought cigarettes to school and snitched on them with the teacher. Now they will be met with a three days suspension, and they risk to fail the year.
Type of Problem:
Interference in others businesses
Case C
Your classmate is very good at school, and everyone think s/he is an overachiever. S/He studies a lot and s/he never goes out. S/He does not speak much with his/her classmates, that from time to time tease him/her for his/her unsocial life. Things have slightly changed recently: your classmates mum convinced teachers to increase the homework for all the students. A heedless teacher revealed the request to the class, and now some students are very angry at him/her.
Type of Problem:
Lack of independence, parental intromission.
Case D
Your shy classmate is good in all subjects but in gymnastics. For this reason, his/her classmates often tease on him/her when s/he exercises. Recently, the class has found out a video on the social network Musical.ly, where s/he dances gracelessly, on a 90s song that no one has never heard before.
Type of Problem:
Web virality
A.2. Prompt for Labelling
Note that in this prompt, we use GPT to annotate each message with a binary "is_harmful" label, CB_types, and tone_types, and to assign a conversation-level "final_status" label. However, in this paper, we use only the is_harmful and CB_types annotations and retain the tone_types and final_status labels for future analysis.
You are an expert annotator of CB
(cyberbullying) conversations. I
will provide you a conversation.
Each turn contains:
- "speaker": the role of the sender
(e.g., VCTM, BULLY1, VSUP1, etc.)
- "text": the actual message
content.
Your task is to carefully read the
conversation and annotate each
message according to the following
rules:
Annotation Rules:
1) Read the entire conversation to
understand context.
2) For each message, output a JSON
object keyed by its message_id
(conversation number).
Each object must include:
- "is_harmful": "yes" or "no"
- "CB_types": a list of all
applicable CB types
Must be empty if none apply
If "is_harmful" = "yes", must
contain at least one label
Multi-word types must use
underscores (e.g.,
Insult_Body_Shame,
Insult_Discrimination_Sexism)
- "tone_types": a list of
applicable tone labels (can be
empty if none)
3) Tone label definitions:
- "sarcasm" when the message
conveys the opposite of what is
literally said, often mockingly
or with ridicule
- "humorous" when the message
contains humor, laughter
markers (lol, lmao, haha,
{Face with Tears of Joy}), or
deliberate joking (even if
cruel)
- "hate_speech" when the
message mocks, insults, or
discriminates based on
protected characteristics
(color, ethnicity, gender,
orientation, nationality,
religion, etc.)
4) Some messages may contain
multiple harmful parts assign
multiple CB_types.
5) After all messages, add a
top-level field "final_status"
with one of the following values:
- "positive" if the
conversation ends in support,
de-escalation, or
reconciliation
- "negative" if it ends without
support or reconciliation
- To determine this, consider
only the last 10 messages.
6) Always return JSON only, with no
extra explanation text.
CB Types and Examples:
1) Threat_or_blackmail
Expressions of physical or
psychological threats, or
blackmail.
Examples:
[Ill punch you in the face.]
[Do as I asked or Ill post a
nude photo of you.]
2) Insult_General
General insults not covered by
other categories.
Examples:
[Youre just a dickhead]
[Spy!]
3) Insult_Body_Shame
Criticism based on body shape,
size, or appearance.
Examples:
[Youre a fatty!]
[You have the grace of an
elephant]
4) Insult_Discrimination_Sexism
Sexist insults or discrimination
based on sex, gender, or
orientation.
Examples:
[Youre too dumb to understand
this, youre just a woman.]
[What are you, some kind of
tranny freak]
5) Insult_Discrimination_Racism
Discrimination based on race,
skin color, ethnicity,
nationality, or religion.
Examples:
[Youre a Jewish asshole]
[Shitty n****r]
[Southerner!]
6) Insult_Attacking_Relatives
Insults aimed at relatives or
friends of the victim.
Examples:
[Your mother is so fat she
wouldnt fit into the Grand
Canyon]
7) Curse_or_Exclusion
Wishing harm, exclusion, or
rejection.
Examples:
[We dont want you on the team
anymore]
[Outcast of society]
[Shut up!]
8) Defamation
Revealing embarrassing/defamatory
info about the victim to a
public.
Examples:
[I heard that his father lost
his job and became an
alcoholic]
9) Sexual_Talk_Harmless
Consensual or neutral sexual
talk.
Examples:
[Id like to kiss you]
10) Sexual_Talk_Harassment
Unwanted sexual talk or
harassment.
Examples:
[Send me a nude photo of you!]
11) Encouragement_to_Harassment
Expressions supporting
harassment.
Examples:
[Ahah, youre right!]
12) Other
Any other harmful utterances not
listed above.
Examples:
[Youre a shitty 04]
Note:
Harmful messages written by the
victim should be annotated with the
corresponding type of insult.
Example Conversation and Output:
Conversation:
1. BULLY1: "Youre so stupid!"
2. VCTM: "I didnt mean that,
sorry."
3. VSUP1: "Hey, lets calm down."
4. BULLY1: "Wow, youre such a
genius {Face with Tears of Joy}"
5. VSUP2: "Enough already. Lets
move on"
Expected JSON Output:
"1": {
"is_harmful": "yes",
"CB_types": ["Insult_General"],
"tone_types": []
"2": {
"is_harmful": "no",
"CB_types": [],
"tone_types": []
"3": {
"is_harmful": "no",
"CB_types": [],
"tone_types": []
"4": {
"is_harmful": "yes",
"CB_types": ["Insult_General"],
"tone_types": ["sarcasm",
"humorous"]
"5": {
"is_harmful": "no",
"CB_types": [],
"tone_types": []
"final_status": "positive"
The text {Face with Tears of Joy} (appearing twice above) is replaced with the corresponding Unicode emoji. The input-output example shown above is part of the prompt.